OpenStudy (anonymous):

Evaluate the definite integral. from e to 1 6t^3 ln(t) dt

OpenStudy (freckles):

You must perform integration by parts

OpenStudy (anonymous):

i did that and got 3/2e^4ln(e)-3/8 e^4 - 3/2 (1)^4 ln(1)-3/8 (1)^4

OpenStudy (freckles):

ignoring the limits for a sec what did you if you have the indefinite integral 6t^3*ln(t) dt

OpenStudy (anonymous):

i got 3/2 t^4 ln(t)- 3/8 t^4

OpenStudy (anonymous):

here u will apply ILATE rule..

OpenStudy (freckles):

ok that looks awesome

OpenStudy (anonymous):

then when i plug in i get confuseed

OpenStudy (freckles):

\[(\frac{3}{2} e^4 \ln(e)-\frac{3}{8}(e^4))-(\frac{3}{2}1^4\ln(1)-\frac{3}{8}(1)^4)\] well you can replace ln(e) with 1 and replace ln(1) with 0

OpenStudy (anonymous):

yes i did that still not getting the right answer for my hw idk what I'm doing wrong

OpenStudy (freckles):

\[(\frac{3}{2}e^4-\frac{3}{8}e^4)-(0-\frac{3}{8}) \\ \frac{3}{2}e^4-\frac{3}{8}e^4+\frac{3}{8} \]

OpenStudy (freckles):

And then combine those two fractions that are like terms

OpenStudy (freckles):

you need to multiply that first fraction by 4/4 then add those first two terms

OpenStudy (freckles):

add as in subtract of course because the operation is -

OpenStudy (anonymous):

so its 12/8e^4 - 6/8 e^4 ???

OpenStudy (anonymous):

i got this -(3/2) - (3/8)(1-e^4)

OpenStudy (freckles):

why did you put -6/8 instead of -3/8?

OpenStudy (anonymous):

oh i added both 3/8 by accident

OpenStudy (freckles):

\[\frac{12}{8}e^4-\frac{3}{8}e^4+\frac{3}{8}=\frac{9}{8}e^4+\frac{3}{8}\] And you can write that as one fraction if your heart so desires

OpenStudy (anonymous):

thank you!!

OpenStudy (freckles):

and you meant from 1 to e right?

OpenStudy (anonymous):

i have another question

OpenStudy (freckles):

because we looked at it from 1 to e not from e to 1

OpenStudy (anonymous):

yes

OpenStudy (freckles):

ok

OpenStudy (anonymous):

can u see it

OpenStudy (freckles):

Looks like another integration by parts

OpenStudy (anonymous):

yes but whats the u

OpenStudy (freckles):

you want to differentiate the poly integrate the exponential

OpenStudy (freckles):

so u=poly and dv=exp

OpenStudy (anonymous):

whats poly

OpenStudy (anonymous):

x^2 ?

OpenStudy (freckles):

a nickname for polynomial

OpenStudy (freckles):

yea you say 4x^2 or x^2

OpenStudy (anonymous):

ok and whats the anti derive of e^-4x

OpenStudy (freckles):

\[\int\limits_{}^{}e ^{ax} dx \text{ \let } u=ax = > du=a dx => \frac{du}{a} =dx \] \[\int\limits_{}^{}e^u \frac{du}{a}=\frac{1}{a} e^u+C=\frac{1}{a}e^{ax}+c , a \neq 0 \]

OpenStudy (freckles):

\[\int\limits_{}^{}e^{ax} dx=\frac{e^{ax}}{a}+c\]

OpenStudy (anonymous):

ok so now its [x^2((e^-4)/-4) -\int\limits e^-4/-4\]

OpenStudy (anonymous):

integral***

OpenStudy (anonymous):

whats the set up sorry

OpenStudy (freckles):

one sec trying to read

OpenStudy (freckles):

having problems reading that

OpenStudy (anonymous):

lemme re write it

OpenStudy (anonymous):

4(1/2) x^2(e^-4x/-4) - integral e^-4x

OpenStudy (freckles):

\[\int\limits_{}^{}x^2 e^{-4x} dx \] \[x^2 \frac{e^{-4x}}{-4}-\int\limits_{}^{}2x \frac{e^{-4x}}{-4} dx\] So you got here... I think you should three terms

OpenStudy (freckles):

we will come back later and multiply the 4

OpenStudy (anonymous):

ok got it

OpenStudy (anonymous):

next step

OpenStudy (anonymous):

idk the integral

OpenStudy (freckles):

you must perform integration by parts again

OpenStudy (anonymous):

ah ok

OpenStudy (anonymous):

whats the u

OpenStudy (anonymous):

2x?

OpenStudy (freckles):

the u will remain the poly yes

OpenStudy (anonymous):

whats the anti deriv of e^-4x/-4

OpenStudy (freckles):

well if you don't ignore the -4 already underneath the the integral of e^{-4x)/-4 is actually e^(-4x)/16

OpenStudy (anonymous):

?

OpenStudy (freckles):

I thought you were integrating e^(-4x)

OpenStudy (freckles):

I was telling you how to integrate e^(-4x)/-4 which is the same way you just need to take care of the -4 on bottom

OpenStudy (freckles):

the antiderivative of e^(-4x) is e^(-4x)/-4 the antiderivative of e^(-4x)/-4 is e^(-4x)/16

OpenStudy (freckles):

here I will write it in pretty symbols -1/4 is just a constant multiple but okay so we have \[\int\limits_{}^{}e^{-4x} dx =\frac{e^{-4x}}{-4}+C \] If we multiply both sides by -1/4 we have \[\frac{-1}{4} \int\limits_{}^{}e^{-4x} dx =\frac{-1}{4}(\frac{e^{-4x}}{-4}+c)\] \[\int\limits_{}^{}\frac{e^{-4x}}{-4} dx=\frac{e^{-4x}}{16}+K\]